# Generative 3D Part Assembly via Part-Whole-Hierarchy Message Passing

### v1

@@ 2 Related Work
Numerous prior studies have approached the challenging 3D part assembly through the joint estimation of part poses. The influential work by [7] introduces an intelligent scissoring technique to tackle this problem for part components. Subsequent research by [2, 12, 10] employs graphical models to capture the semantic and geometric relationships among shape components, enabling exploration in assembly-based shape modeling. PAGENet [14] presents a network that is aware of individual parts and generates semantic parts while estimating their joint poses, while [31] proposes to decompose shapes using a library of 3D parts provided by the user. However, these studies either assumed prior knowledge of part semantics or relied on existing shape databases. In a more practical setting, the authors in [33, 21, 34, 3] focus on the pose estimation of individual parts without relying on shape databases or known semantic information. Specifically, DGL [33] employs a dynamic part graph to iteratively refine the poses of individual parts. A progressive strategy using the recurrent graph learning framework has been investigated in [21]. Additionally, the authors in [34, 3] utilizes the Transformer [28] to model the structural relationships and performs the simultaneous assembly of all parts. We follow this line of work and propose novel improvement to generate structurally-coherent part assemblies.
@@ 3.2 Super-Part Relation Reasoning
For super-part pose prediction, we feed the attention-weighted feature $\mathbf{G}^{\prime}_{i}$ into a prediction head containing four fully-connected layers to obtain the pose $\left\{\left(t^{\prime}_{i},r^{\prime}_{i}\right)\right\}_{i=1}^{M}$. Here we apply tanh operation to the translation vector, which restricts the part center offset to the range of $(-1,1)$. Additionally, for simplicity, we predict the quaternion vector $r^{\prime}_{i}=(r^{\prime}_{i0},r^{\prime}_{i1},r^{\prime}_{i2},r^{\prime}_{i3})$ instead of the parameters of the rotation matrix. We then use the Rodrigues formula [23] to obtain the rotation matrix $\mathbf{R}^{\prime}_{i}$ corresponding to $r^{\prime}_{i}$ one by one as follows:
@@ 4.2 Evaluation Metrics
Mean Part/Connectivity Accuracy (mPA/mCA) PA and CA depend on the Chamfer distance threshold $\tau_{p}$ and $\tau_{c}$ for judging whether the assembly is accurate. In order to provide a more comprehensive evaluation of the performance of the model, we average the results under multiple thresholds to get mean part accuracy (mPA) and mean connectivity accuracy (mCA), formally,
@@
QDS Diversity score (DS) [19, 26] evaluates the diversity of the results as $\text{DS}=\frac{1}{N^{2}}\sum\nolimits_{i,j=1}^{N}(d_{c}(\mathbf{P}_{i}^{\ast},\mathbf{P}_{j}^{\ast}))$, where $\mathbf{P}_{i}^{\ast}$ and $\mathbf{P}_{j}^{\ast}$ represent any two assembled shapes. Based on DS, we propose the quality-diversity score (QDS) as below,
@@
WQDS Based on QDS, we further propose the weighted quality-diversity score (WQDS) as below,
@@ 4.5 Ablation Study
Super-Part Encoder Since the super-part encoder is at the core of our part-whole hierarchy message passing network, we conduct an experiment where we remove this module (denoted as w/o Super-Part Enc.). To mitigate the impact of the number of model parameters, we also create another baseline by increasing the number of parameters of the w/o Super-Part Enc. setting (denoted as w/o Super-Part Enc. + Augmented Param.). This involves increasing the number of parameters in the part encoder and the multi-head attention module, ensuring that the overall number of parameters is on the same order of magnitude as the full model setting. As shown in Table 3, the super-part encoder contributes significantly to the final performance. We also observe that augmenting the model parameters without incorporating the super-part encoders actually leads to a performance drop. This finding further highlights the effectiveness of the proposed part-whole hierarchy network for part assembly.
Loss Functions We now evaluate the importance of individual loss terms in Equation (8). As shown in Table 4, removing the part translation loss $\mathcal{L}_{t}$ leads to a significant performance degradation. Similarly, the remaining loss terms, including the rigid rotation loss $\mathcal{L}_{r}$ and the shape assembly loss $\mathcal{L}_{s}$, demonstrate their significance in facilitating precise part rotations and ensuring the overall coherence of the assembled parts, respectively. This shows that each loss term does play an indispensable role in achieving high-quality assemblies.

### v3

@@ 2 Related Work
Numerous prior studies have approached the challenging 3D part assembly through the joint estimation of part poses. The influential work by [7] introduces an intelligent scissoring technique to tackle this problem for part components. Subsequent research by [2, 12, 10] employs graphical models to capture the semantic and geometric relationships among shape components, enabling exploration in assembly-based shape modeling. PAGENet [14] presents a network that is aware of individual parts and generates semantic parts along with their poses. [31] proposes to decompose shapes using a library of 3D parts provided by the user. However, these studies either assumed prior knowledge of part semantics or relied on existing shape databases. In a more practical setting, the authors in [33, 21, 34, 3] focus on the pose estimation of individual parts without relying on shape databases or known semantic information. Specifically, DGL [33] employs a dynamic part graph to iteratively refine the poses of individual parts. A progressive strategy using the recurrent graph learning framework has been investigated in [21]. Additionally, the authors in [34, 3] utilizes the Transformer [28] to model the structural relationships and performs the simultaneous assembly of all parts. We follow this line of work and propose novel improvement to generate structurally-coherent part assemblies.
@@ 3.2 Super-Part Relation Reasoning
For super-part pose prediction, we feed the attention-weighted feature $\mathbf{G}^{\prime}_{i}$ into a prediction head containing 4 fully-connected layers to obtain the pose $\left\{\left(t^{\prime}_{i},r^{\prime}_{i}\right)\right\}_{i=1}^{M}$. We apply tanh operation to the translation vector, which restricts the range of the part center offset as $(-1,1)$. Additionally, for simplicity, we predict the quaternion vector $r^{\prime}_{i}=(r^{\prime}_{i0},r^{\prime}_{i1},r^{\prime}_{i2},r^{\prime}_{i3})$ instead of the rotation matrix. We then use the Rodrigues formula [23] to obtain the rotation matrix $\mathbf{R}^{\prime}_{i}$ corresponding to $r^{\prime}_{i}$ one by one as follows:
@@ 4.2 Evaluation Metrics
Mean Part/Connectivity Accuracy (mPA/mCA).
PA and CA depend on the Chamfer distance threshold $\tau_{p}$ and $\tau_{c}$ for judging whether the assembly is accurate. In order to provide a more comprehensive evaluation of the performance of the model, we average the results under multiple thresholds to get mean part accuracy (mPA) and mean connectivity accuracy (mCA), formally,
@@
QDS.
Diversity score (DS) [19, 26] evaluates the diversity of the results as $\text{DS}=\frac{1}{N^{2}}\sum\nolimits_{i,j=1}^{N}(d_{c}(\mathbf{P}_{i}^{\ast},\mathbf{P}_{j}^{\ast}))$, where $\mathbf{P}_{i}^{\ast}$ and $\mathbf{P}_{j}^{\ast}$ represent any two assembled shapes. Based on DS, we propose the quality-diversity score (QDS) as below,
@@
WQDS.
Based on QDS, we further propose the weighted quality-diversity score (WQDS) as below,
@@ 4.5 Ablation Study
Super-Part Encoder.
Since the super-part encoder is at the core of our part-whole hierarchy message passing network, we conduct an experiment where we remove this module (denoted as w/o Super-Part Enc.). To mitigate the impact of the number of model parameters, we also create another baseline by increasing the number of parameters of the w/o Super-Part Enc. setting (denoted as w/o Super-Part Enc. + Augmented Param.). This involves increasing the number of parameters in the part encoder and the multi-head attention module, ensuring that the overall number of parameters is on the same order of magnitude as the full model setting. As shown in Table 3, the super-part encoder contributes significantly to the final performance. We also observe that augmenting the model parameters without incorporating the super-part encoders actually leads to a performance drop. This finding further highlights the effectiveness of the proposed part-whole hierarchy network for part assembly.
Loss Functions.
We now evaluate the importance of individual loss terms in Equation (6). As shown in Table 4, removing the part translation loss $\mathcal{L}_{t}$ leads to a significant performance degradation. Similarly, the remaining loss terms, including the rigid rotation loss $\mathcal{L}_{r}$ and the shape assembly loss $\mathcal{L}_{s}$, demonstrate their significance in facilitating precise part rotations and ensuring the overall coherence of the assembled parts, respectively. This shows that each loss term does play an indispensable role in achieving high-quality assemblies.
@@ 5 Conclusion
Acknowledgments
This work was funded, in part, by NSERC DG Grants (No. RGPIN-2022-04636), the Vector Institute for AI, and Canada CIFAR AI Chair. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant.